One pipe can fill a pool 1.25 times faster than a second pipe. When both pipes are opened , they fill the pool in 5 hours. How long would it take to fill the pool if only the slower pipe is used?
Answer by jorel1380(3719)

1/1.25=.8
1/t+1/.8t=1/5
5+6.25=t
11.25=t
.8t=9
The slower pipe would take 11.25 hours alone to fill the pool..